5.) A 14 ml solution of KOH has a concentration of 0.315 M. Calculate the pH and then calculate the amount of KOH in grams in solution.
1
Expert's answer
2020-05-07T14:15:25-0400

1) C(KOH) = n(KOH)/V(KOH) = 0.315/0.014 =22.5 M.

pOH = -Log[OH-] = -Log(22.5) = -1.352

pH = 14 - pOH = 14 - (-1.352) = 15.352

2) n(KOH) = V(KOH)*C(KOH) = 0.315*0.014 = 0.00441 mol

m(KOH) = n(KOH)*M(KOH) = 0.00441*56 = 0.247 g